OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 3(x + 8)^2(x - 8)^3

OpenStudy (campbell_st):

well there are only 2 zeros when x + 8= 0 and x - 8 = 0 the multiplicity of each is determined by the power just think \[(x + 8)^2 = (x + 8)(x + 8)\] so how many times does x + 8 occur.... thats the multiplicity... same reasoning for x - 8 hope this helps.